27/08/2020 – VoIP MSO [Resolved]

We are currently experiencing an issue with VoIP calls across our network, engineers are working on this and will provide an update asap.

-Update- 27/08/2020 09:48
The issue has been identified with one of our up stream carriers who are currently working on an issue and will provide updates shortly.

-Update- 27/08/2020 09:56 Telstra LHC Outage

All powered equipment in Telstra LHC Data Center went offline at 09:17 BST. We have been informed there are multiple fire engines on site and a suspected UPS fire on the 3rd floor where our comms equipment is located. It seems most likely the fire brigade have ordered building power down as part of electrical safety procedure.

As far as we are aware, this affects all customers and carriers within LHC, and we have confirmation that other carriers with passive connections extending outside the building are also showing offline. This is therefore affecting customers who are directly terminated with active circuits at this location. All passive fibre connections remain unaffected, including those passing through the building.

Updates to follow as they arrive from the DC. We sincerely apologise for any inconvenience this may cause.

-Update- 27/08/2020 11:10

Correspondence from LHC DC:
Due to a localised fire in the LHC, we have lost the Green system. This provides power to the lower half of the building. The fire has tripped the breakers supporting the busbar. Engineers are on-site and are working to restore power via generator as we speak.

-Update- 27/08/2020 13:16

We have been made aware by the DC that the services are starting to restore. We are monitoring this carefully and will provide you with an update as soon as we have more information.

-Update- 27/08/2020 14:08

We are seeing call flows return to normal levels but have yet to hear back from the DC and/or Carrier. We will continue to monitor and provide updates as they become available.

-Update- 27/08/2020 16:07

Engineers are still working to improve stability in the network and restore the remaining services we hope this will be complete with in the next 30 minutes. Most customers should have service restored.

-Update- 27/08/2020 16:40

We can see normal call flow across the network but have still yet to get a clear message from the carrier.

-Update- 27/08/2020 17:45

The carrier reports all services are now restored.

-Update- 28/08/2020 9:30

Services are now fully restored, a full RFO will be posted once made available.

06/06/2019 10:00 – Broadband Disruption

We are currently seeing packet loss on many broadband connections, this is currently being investigated and updates will be provided shortly.

UPDATE 01 – 10:45

We have located a back-hull link within our network between our core in Horsham (HOR) and London (THN) that is operating with packet loss at a low level which is affecting services provided by our Horsham facility. This has been removed from service and traffic is flowing via alternative routes.

UPDATE 02 – 11:00

The link removal has restored full service to platforms that operated in part via this link. Investigations are taking place with our fibre backhull provider to ascertain where the fault is. However we suspect this is a common fault further upstream.

UPDATE 03 – 11:29

We are aware some Horsham based broadband services are still seeing issues. Investigations show this to be based around a common fibre link and provider as per our backhull link and investigations have been escalated.

UPDATE 04 – 12:55

We have chased for an update with our suppler who have advised they are seeing issues within there core network affecting other exchanges and services. We have pushed for an escalation due to the severity of the issue.

We apologise for the continued disruption

UPDATE 05 – 14:10

We have been advised the issue is still ongoing and we expect an update by 15:00

UPDATE 06 – 16:01

We have placed further escalations to seniors managers at our suppler due to the length of time this has been ongoing.

UPDATE 07 – 17:00

Our investigations have concluded after speaking with other wholesale providers that this is a result of a Vodafone wholesale issue and a problem within there core network that they are trying to isolate.

UPDATE 08 – 18:10

We have observed packet loss returning to 0% on affected services. This is been confirmed by our supplier however we are continuing to monitor.

UPDATE 09 – 18:50

We have continued to see a good service. Other ISP using Vodafone backhull have advised the same, however we are containing to monitor.

30/05/2019 06:00 – Broadband Disruption

-Update 08:50- We are starting to see some connections come back into service, we are continuing to monitor the situation and will provide more updates shortly.


We are currently aware of an issue effecting lease line and broadband customers. Some services are currently down, we are working with suppliers to get this resolved as soon as possible and will provide updates as we get them.

Apologies for any inconvenience caused.

11/10/2017 – DSL Performance Issues

We are currently aware of an issue effecting DSL connections, we are looking into this and will update shortly.

-Update: 13:24-
We are starting to see connections return back to normal speeds, we are continuing to monitor and will update you shortly.

-Update: 13:12-
Engineers are still working on the issue. We will update with more information as it becomes available.

EasyBOND – Connectivity Issue 16/03/2017

We are currently looking into an issue with the bonded connectivity, we will update with more information shortly.

Update: 10:58
Connections are starting to be reconnect, we are continuing to monitor and will update shortly.

Update: 14:41
Engineers found a “stale” routing issue on C2 that was not renewed until a BGP filter was rebuilt and reset. Traffic routing from C1 was unaffected.

LON01 – EasyIPT – 10/11/2016 – 13:30 *Call Quality*

We are currently aware of an issue effecting call quality across our VoIP Network. Engineers are currently working on this and will provide updates as soon as possible.

UPDATE 13:55 10/11/2016
Call quality has returned to normal standards, engineers are still working to confirm the issue is fully clear and will provide updates as soon as possible.

UPDATE 14:35 10/11/2016
We have now been made aware that a specific Tier One Carrier is currently reporting a major service outage which is affecting customers nationwide at the present time, we are in communication with the supplier in question and will report any additional information to you as and when it is received.

We apologise for any inconvenience this may be causing.

Virgin Media Based Circuits – 03/08/2016 – Incident – Cleared

Update: 15:52
Virgin Media have advised that all repair work is now complete and all services restored. Circuits should no longer be considered at risk.

We apologise the inconvenience caused during the course of this incident.

————-

We are currently aware of an issue affecting a number of on-net Virgin Media based circuits. We believe this is due to a fibre break within the Virgin Media network, no ETR has currently been given.

Services with backup media such as DSL will have been re-routed.

Update: 09:35
Virgin Media have advised that the location of the break has been identified as 7.5km away from a local Virgin Media hub. Splicing engineers are currently attending the closest joint to determine the exact location of the break.

Update: 10:26
Following our previous post, several engineers are now on site identifying the exact location of the fibre break. We are expecting more specific details on the break in due course.

We will keep in contact with the relevant escalation points within Virgin Media to obtain the latest information, as it becomes available.

Update: 11:30
Virgin Media have advised engineers have attended the next joint along from the break and completed Optical time-domain reflectometer (OTDR) testing between the joints. Red light testing is currently underway to confirm the precise location of the break. Virgin Media have also confirmed splicing teams are on site to begin work as soon as the break location is confirmed.

We will post updates as soon as they become available.

Update: 12:58
Virgin Media engineers have located the break between Corbet Place and Brick Lane in London. Engineers are currently accessing the pits this cable passes through to visually inspect the cable and identify if any spare fibres can be used. If any spare fibres are identified they will be prioritised for our services.

Further updates will be posted as soon as available.

Update: 13:22
Virgin Media have advised that spare fibres have been located. Splicing will begin on the first end of the cable shortly and should be complete within 30 minutes. Splicing at the second end of the cable involves more work and is expected to take up to a further 60 minutes to complete.

Further updates will be posted when available.

Update: 14:55
Splicing work on the cables that carry our services has been completed and connectivity restored to all affected circuits. Virgin Media advised work is still ongoing at the site of the break so circuits should still be considered at risk.

We will advise once Virgin Media confirm all repair work complete.

Update: 15:52
Virgin Media have advised that all repair work is now complete and all services restored. Circuits should no longer be considered at risk.

We apologise the inconvenience caused during the course of this incident.

EasyxDSL Cannot Conenct Issues 12/11/2014 – 09:00 *RESOLVED*

We are currently experiencing an issue that is preventing some DSL users connecting to our service and/or constantly reconnecting to our service. Our initial investigations indicate that this is a Entanet issue and we are working with them to resolve as soon as possible. We will update with more information as it becomes available.

— Update 09:17 —
Entanet have identified a hardware failure and advise engineers have been dispatched to replace the faulty hardware. We are still working for a ETA.

— Update – 09:21 —
Any DSL bonded circuits may see reduced capacity but service should remain stable as we use 2 wholesale ISPs with our own interconnects.

— Update – 13:21 —
The issue is still currently on going, Entanet have advised engineers are still en route to replace the hardware. Due to the re routing of traffic you may experience slow speeds and/or packet loss. Please accept our apologies for any inconvenience caused and we will continue to provide updates as they become available.

— Update – 14:01 —
Entanet have advised engineers have arrived on site to replace the faulty hardware.